OpenStudy (anonymous):

One of your friends sends you an email asking you to explain how all of the following expressions have the same answer.

OpenStudy (anonymous):

\[\sqrt[3]{x^3}\] x^1/3 * x^1/3 * x^1/3 \[1 \over x^-1\] \[\sqrt[11]{x^5 * x^4 *x^2}\]

OpenStudy (anonymous):

Compose an email back assisting your friend and highlight the names of the properties of exponents when you use them.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

ahh i composed this mail back to my friend yesterday oly let me see if i can fwd it to you :)

OpenStudy (anonymous):

lol

ganeshie8 (ganeshie8):

here it is, http://openstudy.com/users/rsadhvika#/updates/523961c3e4b0b9d6b2d7707d

ganeshie8 (ganeshie8):

Enjoy..

OpenStudy (anonymous):

thanks ill check it out :)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

need help with the last one

OpenStudy (anonymous):

your other question used addition and mine is multiplcation xD

OpenStudy (anonymous):

\[\sqrt[11]{x^5 * x^4 * x^2}\]

OpenStudy (anonymous):

need to solve that one :)